RAL Progress Log

Thursday 12 Jan 2006

This page is no longer maintained. Please see:

https://wiki.gridpp.ac.uk/wiki/RAL_Tier1_SC3_Log

Thursday 28 July 2005

UKLight connection at risk due to maintenance

Monday 25 July 2005

11:00am

Turned off Lancaster - UKLight router is down for upgrade.

10:38am

Problem with one of the Datastore disk servers over the weekend slowed the rate files were moving to tape

gftp 0441 and 447 rebooted due to lock ups this morning

Friday 22 July 2005

1:21pm

Reduced Lancaster to 1 transfer at a time, seeing lots of Transfer Timeouts in fts logs.

10:32am

Reduced max movers on disk pools to 20, this lead to them all trying to copy 1GB files into the pool on dcache.gridpp, which was still set at 100, set the max movers to 50 on the disk pool, reducded dcache_1 to 20.

All Tier2 transfers have finished, set another 1000 files off for each.

Thursday 21 July 2005

4:15pm

Edinburgh turned themselves off, still having problems

3:10pm

Reactivated Edinburgh, increased them to 4 concurrent transfer

1:20pm

Tape tests have begun, increased stores on each pool to 2.

Wednesday 20 July 2005

5:10pm

Turned off Edinbrugh's channel, they're still having problems with their pool node.

13:56pm

Set up a Gavin's FTS plotting thing. It shows RAL<->T2 transfers. https://fts0344.gridpp.rl.ac.uk:8443/plots/

12:54pm

Admitted defeat - created new database (dcache2) and shifted both srms over to it, Transfers from CERN now working again.

9:33am

Added the Lancaster people to the FTS.

"/C=UK/O=eScience/OU=Lancaster/L=Physics/CN=brian-dteam davies" null
"/C=UK/O=eScience/OU=Bristol/L=IS/CN=tim barrass" null

to

glite-data-transfer-submit-mapfile
glite-data-transfer-query-mapfile
glite-data-transfer-manager-mapfile

1:34am

SRM is still very slow, so slow that CERN's transfers are timing out. srmcp's left for long enough do finish eventually.

csfnfs62 has apparently lost contact. csfnfs62 was the disk server I applied the tcp tuning from the gridfpt servers to...

Edinburgh transfer had moved into Hold status, 32 transfers succeeded, flipped back to Pending/Active, to see how many get done 2nd time through. Imperial had 17 transfers remaining, 311 succeeded, marked all held files as Pending

Tuesday 19 July 2005

1:33pm

Clearing out James Casey's transfers to see if that allows transfers to start from CERN again

Edinburgh have fixed their problem, reactivated their channel

11:15am

Increased Lancaster to 4 files

10:46am

Status:

Lancaster 187 done 766 pending

Imperial 257 done 703 pending

Edinburgh are still inactive

Increased Imperial to 4 transfers

Monday 18 July 2005

9:30pm

Status:

Turned off Edinburgh channel - 175 transfers failed.

Lancaster 40 done 951 pending

Imperial 31 done 963 pending

Edinburgh 0 done 825 pending

7:45pm

dcache.gridpp appears to be suffering under a large number of atlas requests, cancelled outstanding Tier2 transfers to all 3 sites and began new ones going through dcache-tape but still hitting disk pools rather than tape pools.

1:15pm

Started transfer of 1000 files to Edinburgh

10:40am

gftp0442 gridftp log file filled its disk failing transfers, restarted.

Started gridftp door on gftp0440.

Friday 15 July 2005

2:45pm

No lockups seen since Thursday morning, reapplied tuning to gftp0441.

10:50pm

All 4 disk pool logs filled up with : Cell(csfnfs61_3@csfnfs61Domain) : removeFiles : invalid syntax in remove filespec >null<

Emptied log files and restarted pools. Changed periodic file deletion to every 6 hours instead of hourly.

Thursday 14 July 2005

5pm

Transfers with greater than 10 streams failing : Savannah #9552

10:30am

Rebooted gftp0444. Removed tcp tuning parameters from gftp0441 and rebooted it to see if its more stable

Wednesday 13 July 2005

5pm

Rebooted gftp0442

2pm

Rebooted gftp0447

1:15pm

Rebooted gftp0444

1pm

Rebooted gftp0444 and gftp0446

12pm

Rebooted gftp0441 and gftp0445

10am

Rebooted gftp0442

-- Main.rossd - 14 Jul 2005

Edit | Attach | Watch | Print version | History: r23 < r22 < r21 < r20 < r19 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r23 - 2006-01-12 - SteveTraylen
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback